Joe can assemble a computer​ by himself in 60 minutes. Working with an assistant, he can assemble a computer in 50 minutes. How long would it take his assistant to do it working alone?

Answer by Boreal(15235)

in 1 minute, Jan does 1/60 of the job
together, they can do 1/50 of a job in 1 minute
assistant takes x minutes to assemble a computer
assistant does 1/x of the job in a minute
50 minutes to complete it together
(50/60)+(50/x)=1
multiply by 60x
50x+3000=60x
10x=3000
x=300 minutes ANSWER
50/60+50/300=5/6+1/6=1